Suppose you were asked to place square carpet on the floor of a square room. You are told that three times the area of the square room is equal to the product of 6m and its perimeter. How much carpeting materials do you need to cover the entire floor?
Is the answer 64m2 correct?

Answer by ikleyn(52776)

.
The equation, from the condition, is


    3x^2 = 6*(4x).


Simplify


    3x^2 = 24x


    x = 8.


The side length of the square is 8 units; hence, its area is  8^2 = 64 square units.


Your answer is correct.